OpenStudy (anonymous):

Solve: 2(x + 1) = 2x + 2

OpenStudy (anonymous):

2x+2=2x+2 0=0

jhonyy9 (jhonyy9):

2x+2=2x+2 0=0

OpenStudy (anonymous):

i believe theres no solution

OpenStudy (anonymous):

okay but on my test it say 0 no solution 1 all real numbers which one should i choose?